# Pretraining LLM with Latent Thoughts in Continuous Space

### v1

@@ Abstract
The remarkable success of Chain-of-Thought (CoT), which enhances performance by scaling generation steps at test-time, inspires us to ask: can we leverage a similar scaling of computational steps during pretraining to improve the generation of each individual token? To address this, we propose a novel pre-training methodology: Pretraining Language Models with Latent Thoughts. Our approach pretrains a language model (LM) to first generate an intermediate latent thought—the last hidden state of the current position—which is then used as input to predict the actual subsequent token. This additional computational step enables the LM to refine its prediction within unconstrained continuous space. Our experiments demonstrate that, at an identical inference cost, a LM that generates one additional latent thought per token outperforms a standard model with double the parameters. For instance, ours-1.4B (Pythia Arch), pretrained on 300B tokens from the Pile, significantly surpasses the vanilla Pythia-2.8B trained on the same data on both language modeling and a range of general downstream tasks. Furthermore, increasing the number of latent thoughts generated before each actual token—forming a chain analogous to CoT—consistently improves the model’s performance.
Figure 1: Scaling curves comparing our method (Pythia Arch) with the official Pythia suite on the 300B Pile. Our 1.26B model matches the loss of Pythia-2.8B with 55% fewer parameters (left), while our 1.4B model reaches the baseline’s final performance with 62% less training data (right).
@@ 1 Introduction
The conventional wisdom in improving language models—scaling up parameters and data—is facing diminishing returns due to data scarcity (Villalobos et al., 2022; Muennighoff et al., 2023), saturating scaling laws (Hoffmann et al., 2022a; Hackenburg et al., 2025), and prohibitive training overheads (Pati et al., 2023; Narayanan et al., 2021; Li et al., 2024).
@@
This has shifted focus to enhancing model capabilities via test-time scaling (Snell et al., 2024), particularly through methods through Chain-of-Thought (CoT) (Jaech et al., 2024; DeepSeek-AI et al., 2025). CoT achieves remarkable success by generating long reasoning chains for each question, effectively scaling the generation steps and increasing computation per query. While effective, CoT relies on specialized datasets and complex training schemes (Allen-Zhu & Li, 2023; Li et al., 2025; Pang et al., 2025), is confined to a discrete token space, and is ultimately capped by the base model’s capabilities (Yue et al., 2025).
@@
An alternative direction is to scale computation during pretraining. One approach, often termed ”vertical scaling”, deepens the network by reusing parameters (Zeng et al., 2025; Giannou et al., 2023; Geiping et al., 2025; Chen et al., 2025b). However, this can lead to training instabilities (Geiping et al., 2025) and often fails to outperform a standard dense model with a comparable inference budget, limiting its practical utility.
@@
Inspired by the success of CoT in scaling generation steps, we propose a novel ”horizontal scaling” approach: Pretraining Language Models with Latent Thoughts. Instead of deepening the model, our method teaches the LM to scale the generation process for each token. It first generates an intermediate latent thought—the last hidden state of the current position—which is then used as input to predict the actual subsequent token. This allows the model to refine its predictions in an unconstrained continuous space. To maintain training efficiency, we employ the Jacobi iteration (Saad, 2003; Barrett et al., 1994) to parallelize this inherently sequential process.
@@
Our experiments show that, at an identical inference cost, a model trained with one latent thought per token surpasses a standard model with double the parameters. For instance, our 1.4B models, built on Pythia and LLaMA architectures, significantly outperform their vanilla 2.8B counterparts trained on the same data. Our method also proves superior to previous vertical scaling techniques, even when their inference cost is twice as high. Furthermore, increasing the number of latent thoughts to form a chain of latent thoughts—analogous to CoT—before generating each real token consistently improves model performance, further underscoring the potential of our approach.
Figure 2: A comparison between the standard language model and ours. In the standard language model, each token is generated after a single forward pass. In contrast, ours does not immediately sample the output token after one forward pass; instead, it uses the computed last hidden state as the next input embedding for generating the subsequent output token. This allows the language model to think in an unconstrained latent space before producing each token.
2 Related Work
We begin by discussing the two most related works: Coconut (Hao et al., 2024) and PonderLM (Zeng et al., 2025). Coconut finetunes a language model on CoT data, employing a “Chain of Continuous Thought”—represented by the final hidden states—to simulate explicit reasoning steps. This continuous chain is typically applied only after a question is posed. In contrast, our model learns this capability naturally during pretraining on a general corpus, appending a latent token after every token rather than just at the end of a prompt. PonderLM employs a vertical scaling strategy, deepening the model for a single generation step by iteratively re-feeding a “pondering embedding”—a probability-weighted sum of token embeddings—into its input layers. In contrast, our method utilizes a horizontal scaling approach. We extend the generative process for each token by appending latent thoughts, which are directly derived from the last hidden state of the previous computation step. A more comprehensive comparison with related works is provided in Table 1.
Sequential Parameter Sharing to Scale Up Model Depth. This paradigm increases a model’s effective depth by reusing its parameters. Early work like Universal Transformers (Dehghani et al., ) reused entire blocks, while more recent methods refine this by iterating over layers to refine hidden states (Geiping et al., 2025), recycling output states back into the input (Giannou et al., 2023; Saunshi et al., ), or recurrently applying a single layer to critical tokens (Chen et al., 2025a). While these approaches can enhance model capabilities, they often introduce significant inference overhead and training instabilities. In contrast, our horizontal scaling approach avoids the potential instabilities of deep recurrent computations by integrating thought into the sequence length.
Exploring Multiple Solutions through Parallel Computation. This paradigm involves generating multiple candidate solutions in parallel and then selecting the most promising one using a specific criterion. Prominent examples include Best-of-N sampling (Cobbe et al., 2021; Sun et al., 2024; Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024) and Majority Voting (Wang et al., 2022). While effective at improving performance on complex reasoning tasks, these approaches can be computationally inefficient. Furthermore, a key challenge is the difficulty of reliably identifying the best candidate from the generated set, as the verifier or selection heuristic may not be optimal (Stroebl et al., 2024; Hassid et al., 2024).
Scaling Generation Steps. The most prominent method for scaling generation steps is Chain-of-Thought (CoT) (Wei et al., 2022), which elicits reasoning paths from models before they provide a final answer. While effective, this process is often applied at the per-question level. Subsequent work has sought to integrate this ”thinking” process more granularly into generation. One approach involves inserting non-content or ”thinking” tokens into the sequence. For example, Goyal et al. (2023) inserted learnable ”pause” tokens, while others explored discrete planning tokens (Wang et al., 2024) or filler tokens (Pfau et al., 2024). Quiet-STaR (Zelikman et al., 2024) even uses reinforcement learning to generate explicit rationale tokens between output tokens. However, these methods remain constrained to the discrete vocabulary space. In contrast, our work elevates this per-token computation into the continuous latent space.
@@ 2 Related Work
Ours Scaling Generation General Corpus Hidden State Per Token Pretrain
@@ 3 Methodology
In this section, we introduce our pretraining methodology, which trains a language model to first generate an intermediate latent thought before predicting the actual subsequent token. We first establish the notation for a standard Transformer-based language model. Given an input sequence $x=(x_{1},x_{2},\dots,x_{T})$, the model processes the token embeddings $\mathbf{E}_{t}=[\mathbf{e}(x_{1}),\mathbf{e}(x_{2}),\dots,\mathbf{e}(x_{t})]$ using a Transformer architecture. The operation can be formulated as:
$\mathbf{H}_{t}=\text{Transformer}(\mathbf{E}_{t})$
where $\mathbf{e}(\cdot)$ is the token embedding lookup function. The resulting matrix $\mathbf{H}_{t}\in\mathbb{R}^{t\times d}$ contains the sequence of last-layer hidden states. We denote the hidden state at position $t$ as $\mathbf{h}_{t}=\mathbf{H}_{t}[t,:]$.
@@ 3.1 Inference Process
The inference process of our model is straightforward (Figure 2). For each token to be generated, the model first computes its corresponding last hidden state. This hidden state is then used as the input embedding for the subsequent token generation step, mimicking a recurrent thinking process.
@@
3.2 Training Procedure
While inference is sequential, a purely autoregressive training procedure is computationally infeasible for long sequences (e.g., $T=2048$), as it would require thousands of separate forward passes. To address this, we employ the Jacobi iteration to approximate the true autoregressive hidden states, which allows for parallel training (Figure 3). The goal is to find a set of “fixed-point” hidden states $\mathbf{H}^{*}=[\mathbf{h}_{1}^{*},\dots,\mathbf{h}_{T}^{*}]$ that are the output of the model when they are also part of the input. We solve for these states iteratively as follows:
1. Initial Hidden State Estimation (Iteration 0): We begin by performing a single forward pass on the original token embeddings $\mathbf{E}=[\mathbf{e}(x_{1}),\dots,\mathbf{e}(x_{T})]$ to obtain the initial hidden states:
@@ 3.2 Training Procedure
Figure 4: MSE of the last hidden states before and after the ith iteration. The model is the vanilla Pythia-1B tested with 4*2048 tokens.
2. Parallel State Update via Jacobi Iteration (Iteration $k\to k+1$): For each subsequent iteration $k$, we construct a new input sequence by interleaving the original token embeddings with the hidden states from the previous iteration, $\mathbf{H}^{k}$:
@@
3. Loss Computation: After $K$ Jacobi iterations, we form the final input sequence $\mathbf{S}^{K}=[\mathbf{e}(x_{1}),\mathbf{h}_{1}^{K},\dots,\mathbf{e}(x_{T}),\mathbf{h}_{T}^{K}]$. The language modeling objective is then optimized by computing the cross-entropy loss ($L_{1},L_{2},\dots,L_{T}$) at the positions corresponding to the final hidden state inputs. Specifically, the loss $L_{i}$ is computed for predicting the token $x_{i+1}$ from the hidden state $\mathbf{h}_{i}^{K}$. To prevent overfitting to a fixed number of steps, we randomly sample $K$ from $\{2,3,4\}$ for each training instance.
@@
3.3 Position Embedding
When a last hidden state is fed back into the model as an input, it inherits the same positional encoding as its corresponding original token embedding. For instance, the positional encoding for $\mathbf{h}_{i}^{k}$ is identical to that of $\mathbf{e}(x_{i})$ for all iterations $k$.
@@ 4 Experiments
Our evaluation comprises six main components:
  1. 1.
    We first present the results of large-scale pretraining for our model on the 300B-token Pile dataset (Gao et al., 2020), analyzing its scaling behavior and language modeling performance in comparison to the official Pythia suite (Biderman et al., 2023).
  2. 2.
    Next, we assess our model on a broad range of downstream tasks, including general benchmarks and instruction-following, and compare its performance with the official Pythia suite, the PonderLM-Pythia (Zeng et al., 2025), and established models such as OPT (Zhang et al., 2022), Bloom (Le Scao et al., 2023), and TinyLLaMA (Zhang et al., 2024).
  3. 3.
    We further benchmark our approach against several competitive baselines, including Looped Transformer (Giannou et al., 2023), Pause Token (Goyal et al., 2023), PonderLM, and models with doubled parameter counts.
  4. 4.
    To validate the effectiveness of our method on off-the-shelf foundation models, we perform continual pretraining on Llama-3-3B (Grattafiori et al., 2024).
  5. 5.
    Finally, we conduct an ablation study to analyze the impact of key hyperparameters.
@@ 4.1 Large-Scale Pretraining on Pile
We begin by validating our method at scale. We select the Pile (Gao et al., 2020), a substantial 300B-token dataset, as it provides a comprehensive pretraining corpus while remaining computationally tractable.We pretrain models based on the Pythia architecture (Biderman et al., 2023) for two key reasons.. First, its training protocol, including all hyperparameters, is publicly available, enabling a highly controlled experiment where the gains from our method can be isolated. Second, this foundation allows for a direct and rigorous comparison against both the official Pythia models and relevant prior work like PonderLM-Pythia (Zeng et al., 2025).
[…]

### v3

# PonderLM-2: Pretraining LLM with Latent Thoughts in Continuous Space

@@ Abstract
The remarkable success of Chain-of-Thought (CoT), which enhances performance by scaling generation steps at test-time, inspires us to ask: can we leverage a similar scaling of computational steps during pretraining to improve the generation of each individual token? To address this, we propose a novel pre-training methodology: Pretraining Language Models with Latent Thoughts (PonderLM-2). Our approach pretrains a language model (LM) to first generate an intermediate latent thought—the last hidden state of the current position—which is then used as input to predict the actual subsequent token. This additional computational step enables the LM to refine its prediction within unconstrained continuous space. Our experiments demonstrate that, at an identical inference cost, a LM that generates one additional latent thought per token outperforms a standard model with double the parameters. For instance, our PonderLM-2-Pythia-1.4B, pretrained on 300B tokens from the Pile, significantly surpasses the vanilla Pythia-2.8B trained on the same data on both language modeling and a range of general downstream tasks. Furthermore, increasing the number of latent thoughts generated before each actual token—forming a chain analogous to CoT—consistently improves the model’s performance.
Figure 1: Scaling curves comparing PonderLM-2-Pythia with the official Pythia suite on the 300B Pile. Our 1.26B model matches Pythia-2.8B with 55% fewer parameters (left), and our 1.4B model matches Pythia-1.4B with 62% fewer training tokens (middle). On 9 downstream tasks, PonderLM-2-Pythia attains higher average accuracy than TinyLlama at the same training FLOPs (right).
@@ 1 Introduction
The conventional recipe for improving language models—scaling parameters and data—is yielding diminishing returns due to data scarcity (villalobos2022will; muennighoff2023scaling), saturating scaling laws (hoffmann2022training; hackenburg2025scaling), and prohibitive communication overheads (pati2023computation; li2024understanding).
@@
This has shifted focus towards enhancing model capabilities via test-time scaling (snell2024scaling), particularly through methods based on Chain-of-Thought (CoT) (jaech2024openai; deepseekai2025deepseekr1incentivizingreasoningcapability). CoT achieves remarkable success by generating long reasoning chains for each question, effectively scaling the generation steps and increasing computation per query. While effective, CoT relies on specialized datasets and complex training schemes (allen2023physics; li2025llms; pang2025bolt), is confined to a discrete token space, and is ultimately capped by the base model’s capabilities (yue2025does).
@@
An alternative direction is to scale computation during pretraining. One approach, often termed "vertical scaling", deepens the network by reusing parameters (zeng2025pretraining; giannou2023looped; geiping2025scaling; chen2025innerthinkingtransformerleveraging). However, this can lead to training instabilities (geiping2025scaling) and often fails to outperform a standard dense model with a comparable inference budget, limiting its practical utility.
@@
Inspired by the success of CoT in scaling generation steps, we propose a novel "horizontal scaling" approach: Pretraining Language Models with Latent Thoughts. Instead of deepening the model, our method teaches the LM to scale the generation process for each token. It first generates an intermediate latent thought—the last hidden state of the current position—which is then used as input to predict the actual subsequent token. This allows the model to refine its predictions in an unconstrained continuous space. To maintain training efficiency, we employ the Jacobi iteration (saad2003imsls; barrett1994templates) to parallelize this inherently sequential process.
@@
Our experiments show that, at an identical inference cost, a model trained with one latent thought per token surpasses a standard model with double the parameters. For instance, our PonderLM-2-1.4B models, built on Pythia and LLaMA architectures, significantly outperform their vanilla 2.8B counterparts trained on the same data. Our method also proves superior to previous vertical scaling techniques, even when their inference cost is twice as high. Furthermore, increasing the number of latent thoughts to form a chain of latent thoughts—analogous to CoT—before generating each real token consistently improves model performance, further underscoring the potential of our approach.
@@
PonderLM-2 Scaling Generation General Corpus Hidden State Per Token Pretrain
2 Related Work
We begin by discussing the two most related works: Coconut (hao2024training) and PonderLM (zeng2025pretraining). Coconut finetunes a language model on CoT data, employing a “Chain of Continuous Thought”—represented by the final hidden states—to simulate explicit reasoning steps. This continuous chain is applied only after a question is posed. In contrast, our model learns this capability naturally during pretraining on a general corpus, appending a latent token after every token rather than just at the end of a prompt. PonderLM employs a vertical scaling strategy, deepening the model for a single generation step by iteratively re-feeding a “pondering embedding”—a probability-weighted sum of token embeddings—into its input layers. In contrast, our method utilizes a horizontal scaling approach. We extend the generative process for each token by appending latent thoughts, which are directly derived from the last hidden state of the previous computation step. A more comprehensive comparison with related works is provided in Table 1.
Sequential Parameter Sharing to Scale Up Model Depth. This paradigm increases a model’s effective depth by reusing parameters. Early work such as Universal Transformers (dehghaniuniversal) reused entire blocks; later methods iterate layers to refine hidden states (geiping2025scaling), feed outputs back as inputs (giannou2023looped; saunshireasoning), or recurrently apply a single layer to critical tokens (chen2025inner). Despite gains, these approaches often incur inference overhead and training instability. In contrast, our horizontal scaling avoids deep recurrent computation by integrating thought into the sequence length.
Exploring Multiple Solutions through Parallel Computation. This paradigm generates multiple candidate solutions in parallel and selects the most promising one via a task-specific criterion. Representative examples include Best-of-$N$ sampling (cobbe2021training; sun2024fast; gui2024bonbon; amini2024variational; sessa2024bond) and Majority Voting (wang2022self). While effective for complex reasoning, these approaches are often computationally inefficient. Moreover, reliably identifying the best candidate remains challenging, since the verifier or selection heuristic may be suboptimal (stroebl2024inference; hassid2024larger).
Scaling Generation Steps. The most prominent method for scaling generation steps is Chain-of-Thought (CoT) (wei2022chain), which elicits reasoning paths from models before they provide a final answer. While effective, this process is often applied at the per-question level. Subsequent work has sought to integrate this "thinking" process more granularly into generation. One approach involves inserting non-content or "thinking" tokens into the sequence. For example, goyal2023think inserted learnable "pause" tokens, while others explored discrete planning tokens (wang2024guidinglanguagemodelreasoning) or filler tokens (pfau2024let). Quiet-STaR (zelikman2024quietstarlanguagemodelsteach) even uses reinforcement learning to generate explicit rationale tokens between output tokens. However, these methods remain constrained to the discrete vocabulary space. In contrast, our work elevates this per-token computation into the continuous latent space.
@@ 3 Methodology
In this section, we introduce our pretraining method, which trains a language model to generate an intermediate latent thought before predicting the next token. For a standard Transformer LM, given an input sequence $x=(x_{1},\dots,x_{T})$, we compute token embeddings $\mathbf{E}_{t}=[\mathbf{e}(x_{1}),\dots,\mathbf{e}(x_{t})]$ and obtain last-layer hidden states
$\mathbf{H}_{t}=\mathrm{Transformer}(\mathbf{E}_{t}),$ (1)
where $\mathbf{e}(\cdot)$ is the embedding lookup. The hidden state at position $t$ is $\mathbf{h}_{t}=\mathbf{H}_{t}[t,:]$.
Figure 2: Comparison between standard LM and PonderLM-2. Left: standard autoregressive decoding, where each token is sampled after a single forward pass on the current prefix. Middle: PonderLM-2 inference, which delays token sampling and feeds the last hidden state back as the next-step input embedding, enabling latent thinking before emitting each token. Right: the resulting sequential dependency in a naive training unrolling. Red arrows indicate an attention dependency: computing $\mathbf{h}_{i}$ requires the prefix state $\mathbf{h}_{i-1}$ via causal attention, yielding $\mathbf{h}_{1}\rightarrow\mathbf{h}_{2}\rightarrow\cdots$ and preventing full parallelization. This motivates our Jacobi-iteration-based parallel training in Figure 3.
@@ 3.1 Inference Process
The inference process of our model is straightforward (Figure 2). For each token to be generated, the model first computes its corresponding last hidden state. This hidden state is then used as the next-step embedding, mimicking a recurrent thinking process.
3.2 Parallel Training via Jacobi Iteration
As illustrated in Figure 2, our inference introduces a recurrent feedback mechanism: after computing the last-layer hidden state at step $i$, we feed it back as the input embedding for generating step i+1i{+}1. This induces an explicit left-to-right dependency (Figure 2, right): due to causal attention, computing $\mathbf{h}_{i}$ depends on the already-computed prefix states $\{\mathbf{h}_{j}\}_{j<i}$, so a naive token-wise unrolling would require $\mathbf{h}_{1}\rightarrow\mathbf{h}_{2}\rightarrow\cdots\rightarrow\mathbf{h}_{T}$, i.e., $T$ sequential forward passes for a length-$T$ sequence. Such a purely sequential training procedure is computationally infeasible for long contexts (e.g., T=2048T{=}2048). To enable efficient parallel training while remaining consistent with sequential inference, we approximate these autoregressive hidden states via Jacobi iteration, which refines all positions in parallel for a small number of rounds (Figure 3). The goal is to find a set of fixed-point hidden states $\mathbf{H}^{*}=[\mathbf{h}_{1}^{*},\dots,\mathbf{h}_{T}^{*}]$ that satisfy a self-consistency condition: they are provided as part of the input and reproduced by the model as outputs. We solve for these states iteratively as follows:
@@ 3.2 Parallel Training via Jacobi Iteration
1. Initial Hidden State Estimation (Iteration 0): We first run a single forward pass on the original token embeddings $\mathbf{E}=[\mathbf{e}(x_{1}),\dots,\mathbf{e}(x_{T})]$ to obtain the initial hidden states:
@@
2. Parallel State Update via Jacobi Iteration (Iteration $k\to k+1$): For each subsequent iteration $k$, we construct a new input sequence by interleaving the original embeddings with the hidden states from the previous iteration, $\mathbf{H}^{k}$:
@@
3. Loss Computation: After $K$ Jacobi iterations, we form the final input sequence $\mathbf{S}^{K}=[\mathbf{e}(x_{1}),\mathbf{h}_{1}^{K},\dots,\mathbf{e}(x_{T}),\mathbf{h}_{T}^{K}]$. The language modeling objective is then optimized by computing the cross-entropy loss ($L_{1},L_{2},\dots,L_{T}$) at the positions corresponding to the final hidden state inputs. Specifically, the loss $L_{i}$ is computed for predicting the token $x_{i+1}$ from the hidden state $\mathbf{h}_{i}^{K}$. To prevent overfitting to a fixed number of steps, we randomly sample $K$ from $\{2,3\}$ for each training instance. We further ablate the effect of this randomness in Appendix E.
@@
3.3 Jacobi Iteration: Convergence and Consistency with Sequential Inference
Our parallel training can be formulated as a fixed-point iteration over hidden states. Given token embeddings $\mathbf{E}=[\mathbf{e}(x_{1}),\dots,\mathbf{e}(x_{T})]\in\mathbb{R}^{T\times d}$, let $\mathbf{H}^{(k)}=[\mathbf{h}^{(k)}_{1},\dots,\mathbf{h}^{(k)}_{T}]\in\mathbb{R}^{T\times d}$ denote the hidden states at Jacobi iteration $k$. In each iteration, we feed the interleaved sequence $\mathbf{S}^{(k)}=[\mathbf{e}(x_{1}),\mathbf{h}^{(k)}_{1},\dots,\mathbf{e}(x_{T}),\mathbf{h}^{(k)}_{T}$] into the same masked Transformer and read out the updated states. Equivalently, this defines an operator $\Phi(\cdot;\mathbf{E})$ such that
$\mathbf{H}^{(k+1)}=\Phi(\mathbf{H}^{(k)};\mathbf{E}),\qquad\mathbf{H}^{*}=\Phi(\mathbf{H}^{*};\mathbf{E}).$
Finite-step stabilization from causality.
By autoregressive causality, the computation of $\mathbf{h}_{i}$ depends only on the prefix $\{1,\dots,i\}$. Hence, once $\{\mathbf{h}_{1},\dots,\mathbf{h}_{i-1}\}$ have stabilized, the update of $\mathbf{h}_{i}$ becomes fixed and will not change afterwards. Therefore, the iteration stabilizes left-to-right and reaches a fixed point in at most $T$ iterations.
Consistency with sequential inference.
Let $\mathbf{H}_{\mathrm{seq}}$ denote the hidden states produced by standard sequential autoregressive inference under the same model. At convergence, both $\mathbf{H}^{*}$ and $\mathbf{H}_{\mathrm{seq}}$ satisfy the same causal update constraints, implying that the parallel objective is consistent with sequential inference. We empirically validate rapid convergence and alignment to $\mathbf{H}_{\mathrm{seq}}$ in Section 4.6.
3.4 Position Embedding and Context Window
When feeding the last hidden state back as input, it preserves the original token’s positional encoding. Thus the model’s usable context window does not shrink (e.g., it is not “cut in half”): each token stays anchored to its original position.
[…]
@@ 4 Experiments
We conduct a comprehensive evaluation to validate both the effectiveness and efficiency of our method. We first scale up training on the 300B-token Pile to study scaling behavior and language modeling performance (Section 4.1). We then assess generalization on downstream benchmarks and instruction-following (Alpaca fine-tuning, MT-Bench) against strong open baselines (Section 4.2). To contextualize efficiency, we further compare with competitive compute-scaling methods (Looped Transformer, Pause Token, PonderLM) and a $2\times$-parameter oracle under matched and higher inference FLOPs budgets (Section 4.3). We also test plug-and-play effectiveness by continual pretraining an off-the-shelf LLaMA-3-3B foundation model (Section 4.4). Finally, we study complementarity with test-time scaling on GSM8K (Section 4.5) and provide convergence diagnostics of the Jacobi iteration (Section 4.6) together with ablations on key hyperparameters (Section 4.7).
@@ 4.1 Large-Scale Pretraining on Pile
We begin by validating our method at scale. We select the Pile (gao2020pile), a substantial 300B-token dataset, as it provides a comprehensive pretraining corpus while remaining computationally tractable. We pretrain models based on the Pythia architecture (biderman2023pythia).